2. A lethal recessive condition is responsible for the death of 1 out of every 10,000 babies born in the United States. p2+2p can someone please help me with the question “c” asap with an explanation please.
Answer #1

Ans - Since you have found

q=0.01

since the equation is of HWE is

p^2 + 2pq + q^2 = 1

now q=0.01 we will find p

Since according to HWE (p+q)^2 =1

so squaring on both sides

p + q = 1

p + 0.01 = 1

p = 1-0.01

p= 0.99

Now if p and q satisfies the HWE then.its is actually in population.

(0.99)^2 + 2×0.99×0.01 + (0.01)^2

on calculating above we get this equals to 1

Hence the above gene is actually present in the HWE in this population.